# Microsoft Research Montréal, Canada

### v1

@@ 2.1.3 Consequence: Marginalization
Mechanism: Appropriation of data labor Li et al. [126] define data labor as “activities that produce digital records useful for capital generation”. The term encompasses both witting labor activities—as in the case of crowdwork [15], peer production [199, 200], and content moderation [81]—and unwitting activities such as user behavior data and other data generated when users interact with and participate on the platforms. Data labor also encompass the generation of artefacts by writers [44, 45], artists [211, 212], and programmers [210] among others outside of the AI development process that are nonetheless extracted from the web and fed in as training data to generative AI models. Appropriation of data labor in this context includes both (i) the uncompensated appropriation of works by writers, authors, programmers, and peer production communities like Wikipedia [44, 45, 212, 210, 182, 33, 34, 215, 213, 17, 135, 41, 40, 80, 216], and (ii) under-compensated crowdwork for data labeling that have been instrumental in the development of these technologies [163, 230, 196, 15, 94, 233, 95]. It is particularly harmful when these technologies developed on appropriated labor is then employed to displace and automate the jobs of those whose labor was already appropriated in the process [16, 51, 215]. Introduction of such automation may involve vicious cycles of skill-transfer from people to AI models whereby every profession is replaced by corresponding lesser-paid gigified equivalent as auditing and editing of model outputs [85], and the AI model capabilities continue to improve by learning from workers’ inputs while the workers themselves progressively lose their economic value and power, get deskilled, and even relegated into the role of moral crumple zones [66]. This is a critical challenge in the context of information access because (i) the devaluation of writers and artists have direct implications for the quality of content on the web, and (ii) these automated content generation tools are starting to get incorporated directly in information access platforms [164]. Similar concerns of commodification and appropriation have also been raised in other information and knowledge access contexts such as in the enterprise [74].
@@
AI for me, data labor for thee. Another pernicious aspect of AI data labor dynamics is how it mirrors and reifies racial capitalism and coloniality, employs global labor exploitation and extractive practices, and reinforces the global north and south divide [92, 24, 154, 117, 49, 147, 194]. These dynamics encompass accruing the benefits of generative AI to privileged western population who speak English or other data-rich languages while data labor is relegated to the global south where these technologies struggle to find relevance. Communities that significantly contribute to AI data labor may even find their own linguistic styles being labeled AI-ese [99] and being forced to repeatedly prove their own humanity [57, 137]. Attempts to bridge the global north-south data gap also in turn may further intensify surveillance capitalism and data extractive practices in the global south [43].
@@
Mechanism: Bias amplification LLMs and other generative models reproduce and amplify harmful biases and stereotypes from their training datasets [29, 36, 83, 28, 22, 11] which can lead to allocative and representational harms [53]. Harms may also materialize from demographic blindness [74] when the model (or, the system it is embedded in) treats different individuals and groups as alike when it in fact is unwarranted. Examples may include the handling of certain languages as one homogeneous entity without regards for sociolects or dialects [27] or holding different perspectives as equally valid without considerations for historical context or structural dynamics of power. These biases are concerning in the context of information access systems that are responsible for supporting informed citizenry and functioning democracies, health literacy, and knowledge production among other societal needs.
@@
Mechanism: AI doxing We introduce the phrase AI doxing to describe the act of leaking private information of people by an AI system. Weidinger et al. [225] note that this may be caused by models leaking private information (e.g., address and telephone number) present in their training data [38] or when these models are employed to predict people’s sensitive attributes (e.g., political and sexual identities) based on what is known about them publicly [120, 156, 169, 235]. Private information in the training data is a challenge even if the datasets have been sourced from the public web because these models may continue to regurgitate that information after it has been removed from the web or these models may bypass safety measures that would prevent such information from surfacing through web search—e.g., the information may be protected by robots.txt that blocks popular search crawlers but misses crawler bots that specifically collect data for AI model training. In many contexts, applications of these models to predict people’s private information may be based on shaky scientific grounds [12, 209], to put it mildly. However, such applications of these models may contribute to serious harms and discrimination regardless of their accuracy as long as some people are convinced of its predictive power and employ it to marginalize others. AI doxing may also take other forms such as reverse-image-search [19], a functionality supported by some search engines, that may be abused for stalking and harassment.
@@ 2.1.4 Consequence: Innovation decay
Mechanism: Industry capture The compute and data moat that concentrates power in hands of big tech, as discussed earlier in §2.1.2, also creates significant barriers to entry for academic research. These barriers limits academic AI research to handful of institutions that have the necessary means and connections to industry who provide access to compute and data resources to incentivize research in areas of their economic interests. Academics who want to contribute to the research on large scale AI systems or critique their sociotechnical impacts are pressured to play well with institutions holding monopolistic control over compute, data, and systems [148]. Access to “open access’ models—without the compute and data necessary to build them from scratch—allows academic researchers to invest in finding more effective applications of these technologies that serve industry interests but never to reimagine them to be radically different. Students and other academics who may someday want to work in industry are shepherded into integrating themselves into this homogenized research agenda.
This industry capture [229] subjugates academic research to the narrow profit-driven sociotechnical imaginaries of big tech and Silicon Valley capitalism [143], and thwarts any research that may not be immediately monetizable or challenges the status quo of power concentration. As Mitra [143] asks: “Whose sociotechnical imaginaries are granted normative status and what myriad of radically alternative futures are we overlooking?” Narratives of the inevitability of these technologies that are hyped up to be both transformative forces for society and simultaneously posing existential risks for humanity (often purported by the same actors) only bolsters their imagined importance to accumulate increasing global investments, including from governments. Researchers who care about sociotechnical impact and ecological sustainability are kept busy enumerating the harms of constantly emerging new AI technologies and chasing potential mitigations instead of imagining and developing systems for social good. Even as the grounded risks from these technologies (such as those discussed here) gather consensus from academic communities and civil society, the option of slowing down their deployment is taken off the table citing “competitive pressures”, or only framed to address imagined future harms [77] in ways that only contributes to further reinforcement of monopolistic powers of those who have already added these technologies to their arsenals. Critical research on sociotechnical harms of AI is also under risk from controversial ideologies [76, 75] that attempts to displace calls for structural changes with altruism and misdirect attention from concerns of real harms to marginalized peoples today to unsubstantiated imagined future concerns. Even the framing of “responsible AI” seems to shift the frame ever so slightly from centering the field on concerns of ethics, accountability, and transparency to one that presupposes that AI can be made responsible and that it is the technology, not the privileged institutions who wield its powers, that should be challenged. This has led some sociotechnical researchers in AI to explicitly draw attention to how these systems shift power (e.g., [110, 28, 141, 74]) and to prioritize research guided by alternative visions for our sociotechnical futures grounded in universal emancipation and social justice [143].
@@
As generative AI starts to accumulate the lion’s share of research investments, it may starve out other areas of information access research. Generative AI has had exciting but limited deployments in information access systems today. There are significant open challenges to making these models broadly useful, including but not limited to concerns of potential sociotechnical harms. There is a risk that if these challenges are not mitigated in spite of the extensive resources already invested on them at present that there may be calls for even larger investments in future prompted by sunk cost fallacy. [^13] https://en.wikipedia.org/wiki/Sunk_cost#Fallacy_effect It would be astute for the IR community to consciously continue to invest on research on systems and applications that our society needs beyond what these emerging AI technologies make plausible [178, 143].
@@
Mechanism: Pollution of research artefacts Risks to academic research from generative AI may also emerge through the applications of these models in IR scholarship—e.g., for authoring scientific papers and peer reviewing. There is evidence that researchers in computational sciences are already leveraging these tools [130], sometimes with hilariously terrible outcomes [161]. While the use of language models for light editing may eventually fall within the norms of socially acceptable behavior in research, their application in scholarship does raise concerns of plagiarism and scientific inaccuracies. This is an area that currently has more questions than answers and the IR community would benefit from proactively considering potential implications of this trend on future IR research.
@@ 2.1.5 Consequence: Ecological impact
Another important consequence of generative AI is its impact on our environment. In this context it is important for us to consider the direct environmental cost of developing and deploying generative AI systems at scale as well as the potential impact of these technologies on the climate change discourse online.
@@
Mechanism: Persuasive advertising Generative AI may not only negatively impact the environment through increasing demand for natural resources and increasing generation of waste, but it may also supercharge climate change disinformation [72, 7, 9, 59, 188, 47]. For example, the fossil-fuel industry may attempt to sway public opinion through advertising that leverages generative AI’s persuasion capabilities discussed in §2.1.2. Persuasive advertising may also be employed by other environmentally-unfriendly business models like fast-fashion [46]. While the direct ecological cost of generative AI justifiably garners lots of attention, its potential impact on the politics of climate change also deserves scrutiny.
@@ 2.2.1 Risks to society
Information access is a critical need of any democratic society and a necessary ingredient for social transformation [100, 166, 82, 48, 84]. It is also a social determinant of economic progress [236, 149] and health [145]. Disruptions to the information ecosystem bears potentially grave risks to most aspects of our social lives. Our world is currently facing a confluence of pandemic [177, 202, 68], rising global conflicts [201, 205], and escalating climate catastrophes [158, 106, 167] that are pushing us towards precarity. Our information ecosystems are already struggling under the weight of misinformation and disinformation that in this critical moment is eroding public trust in online platforms, institutions, and each other. It is imperative that researchers and developers of information access systems prioritize safeguarding social interests and be vigilant in considering potential risks of disruption and ecosystem collapse when integrating generative AI technologies in the IR stack.
The risks to society are not just from potential disruptions of the information ecosystem, but also from how these technologies simultaneously concentrate power away from those at the margins of society. As institutions that develop and operate these technologies are themselves beneficiaries of this concentration, we need democratic oversights over these platforms. If these technologies further exacerbate already worsening wealth and power inequities, it may pose severe threats to democratic institutions and human rights. We must also broaden our frame of discourse beyond how these technologies may make things worse for those at the margins while accruing benefits for those already privileged. We must consider the opportunity cost of not reimagining information access in light of our boldest sociotechnical ambitions of human emancipation, culture, and knowledge production, instead of being constrained solely by what these emerging technologies make plausible and the homogenized visions put forth by elite institutions who wield these technologies [143].
@@ 2.2.2 Risks to IR research
There is a tenable threat to academic IR research from industry capture that may happen through the confluence of different factors including the distancing of academic researchers from the data and compute they need to perform their work and how the prevalent narratives about the inevitability of AI technologies shapes what computational research gets funded. The concentration of access to these technologies in a handful of elite institutions shapes even what is considered “foundational” or even “AI”. Research on generative AI cannot be allowed to be performed only in the context of corporate economic interests while academia is hollowed out and prevented from exploring radical new methods that challenge the status quo. Instead, the IR community must be empowered with both the space and the resources necessary to explore a diversity of these visions and critique dominant narratives. Otherwise, we risk relegating all IR research to just being extensions of industrial system development.
@@ 2.2.3 Risks to environment
Information access provides one of the large scale application settings for generative AI. However, the impact of such wide-scale deployment of these technologies on the impending climate crisis from anthropogenic global warming should be a critical consideration. In contrast to many “existential risks” that have become a subject of popular discourse in recent years but lack adequate scientific basis, the climate impact of generative AI poses a more substantial existential risk for our species. As we discussed in §2.1.5, these concerns include not just the ecological cost of developing and deploying generative AI technologies but also their impact on online discourses around climate change.
@@ 3.1 Evaluating the impact of generative IR applications
Evaluating the impact of generative IR applications requires methods, as do data-informed interventions to steer that impact. Creating a LLM-based demo has become exceedingly easy. Understanding the impact of a system when it gets used in real life contexts, and getting to a high quality experience for a wide variety of users, is much harder. Standards for impact assessment have not kept up a similar pace as tech developments. Klaaf points out the need to carefully consider the differences in value alignment of the goals of a system, and safety considations, harms and risk  [116]. A wide range of online, offline, and human-assisted evaluations are possible -and necessary- to get a full sense of the impact of a system.
@@ 3.2 Threat identification, assessment, and modeling
When the emergence of a new technology or application becomes apparent, the assessment of whether this poses risks or opportunities within specific domains poses a challenge. Before development of a system, threats and opportunities can be identified. As Kapoor et al. [112] point out, it’s crucial to not evaluate the risks and impact of new systems in isolation, but rather in comparison with existing technologies. For example, the impact of usage of foundation models in search should be compared to existing web search. For this purpose, Kapoor et al., present an evaluation framework focus on marginal risks, applied to Open Foundation Models. Their framework is based on threat identification work from cybersecurity and consists of six steps necessary to demonstrate such marginal risk. These steps are: 1) threat identification, 2) evaluating existing risk absent open foundation models, 3) considering existing defenses absent open foundation models, 4) evidence of marginal risk of open foundation models, 5) ease of defending against new risks, and 6) outlining uncertainty and assumptions. Note that this framework does not set exact assessment criteria, but rather defines the steps to get to such evaluations.
In practical settings, this might mean having to select standards for the development process (e.g. emerging standards from organizations such as NIST [8] or ISO [1], company-specific standards such as Microsoft’s Responsible AI Standard v2 General Requirements [2], or following new (local) legal requirements). However, mapping out potential consequences and identifying mechanisms that introduce risks in the specific context of a system needs to go much further. How to disrupt potential negative mechanisms in order to mitigate those risks requires gauging a wide range of consumer-side impacts [64], but also wider societal impacts. That includes framework focused on worker consequences [74], or practical methods focused on reducing the (legal) risks of using certain types of copyrighted or restricted training data vs. expected performance gains  [142].

### v2

@@ 2.1.3 Consequence: Marginalization
Mechanism: Appropriation of data labor Li et al. [129] define data labor as “activities that produce digital records useful for capital generation”. The term encompasses both witting labor activities—as in the case of crowdwork [15], peer production [208, 209], and content moderation [83]—and unwitting activities such as user behavior data and other data generated when users interact with and participate on the platforms. Data labor also encompasses the creation of artefacts by writers [45, 46], artists [220, 221], and programmers [219] etc. outside of the AI development process that are nonetheless extracted from the web and fed in as training data to generative AI models. Appropriation of data labor in this context includes both (i) the uncompensated appropriation of works by writers, authors, programmers, and peer production communities like Wikipedia [45, 46, 221, 219, 191, 34, 35, 224, 222, 17, 141, 42, 41, 82, 225], and (ii) under-compensated crowdwork for data labeling that has been instrumental in the development of these technologies [170, 239, 205, 15, 96, 242, 97].
It is particularly harmful when technology developed on appropriated labor is then employed to displace and automate the jobs of those whose labor was appropriated [16, 52, 224].Introduction of such automation may involve vicious cycles of perceived skill-transfer from people to AI models whereby professional jobs are replaced by corresponding lesser-paid gigified equivalent as auditing and editing of model outputs only [87]. Proprietary AI model capabilities may then continue to improve by learning from workers’ inputs, while workers progressively lose their economic value and power, or are even relegated into the role of moral crumple zones [68].
This is a critical challenge in the context of information access because (i) the devaluation of writers and artists have direct implications for the quality of content on the web, and (ii) these automated content generation tools are starting to get incorporated directly in information access platforms [171]. Similar concerns of commodification and appropriation have also been raised in other information and knowledge access contexts such as in the enterprise [76].
@@
AI for me, data labor for thee. Another pernicious aspect of AI data labor dynamics discussed in the literature is how they can mirror and reify racial capitalism and coloniality, employ global labor exploitation and extractive practices, and reinforce the global north and south divide [94, 25, 161, 120, 50, 154, 203]. While worldwide jobs might be created in certain cases, the workers are typically low-paid and deprived of any share of the profit made from technologies built with their labor.These dynamics encompass accruing the benefits of generative AI to privileged populations, while data labor is relegated to already marginalized populations, for example in the global south. Communities that significantly contribute to AI data labor may even find their own linguistic styles being labeled AI-ese [101] and being forced to repeatedly prove their own humanity [58, 143]. Attempts to bridge the global north-south data gap also in turn may further intensify data extractive practices in the global south [44].
@@
Mechanism: Bias amplification LLMs and other generative models reproduce and amplify harmful biases and stereotypes from their training datasets [30, 37, 85, 29, 23, 11] which can lead to allocative and representational harms [54]. Harms may also materialize from demographic blindness [76] when the model (or the system it is embedded in) treats different individuals and groups as alike when, in fact, it is unwarranted. Examples may include the handling of certain languages as one homogeneous entity without regards for sociolects or dialects [28] or holding different perspectives as equally valid without considerations for historical context or structural dynamics of power. These biases are concerning in the context of information access systems that are responsible for supporting informed citizenry and functioning democracies, health literacy, and knowledge production among other societal needs.
@@
Mechanism: AI exploitation and doxing “AI doxing” can describe the act of leaking people’s private information by an AI system. Weidinger et al. [234] note that this may be caused by models leaking private information (e.g., address and telephone number) present in their training data [39] or when these models are employed to predict people’s sensitive attributes (e.g., political and sexual identities) based on what is known about them publicly [123, 163, 177, 244]. Private information in the training data is a challenge even if the datasets have been sourced from the public web because models may continue to regurgitate that information after it has been removed from the web, or bypass safety measures that would prevent such information from surfacing through web search—e.g., the information may be protected by robots.txt that blocks popular search crawlers but misses crawler bots that specifically collect data for AI model training. In many contexts, applications of these models to predict people’s private information may be based on shaky scientific grounds [12, 218], to put it mildly. However, such applications may still contribute to serious harms and discrimination regardless of their accuracy as long as some people are convinced of their predictive power and employ them to marginalize others. AI doxing may also take other forms such as reverse-image-search [20], a functionality supported by some search engines, that may be abused for stalking and harassment. In turn, exploitative materials produced with GenAI (such as deep fake revenge porn, or CSAM) might be amplified.
@@ 2.1.4 Consequence: Innovation decay
Mechanism: Industry capture The compute and data moat that concentrates power in the hands of big tech, as discussed earlier in §2.1.2, also creates significant barriers to entry for academic research. These barriers limits academic AI research to a handful of institutions that have the necessary means and connections to industry who provide access to compute and data resources to incentivize research in areas of their economic interests. Academics who want to contribute to research on large scale AI systems or critique their sociotechnical impacts are pressured to play well with institutions holding monopolistic control over compute, data, and systems [155]. Access to “open access” models—without the compute and data necessary to build them from scratch—allows academic researchers to invest in finding more effective applications of these technologies that serve industry interests, but not to reimagine / rearchitect them to in radically different ways. Students and other academics who may someday want to work in industry are shepherded into integrating themselves into this homogenized research agenda.
Such “industry capture” [238] allows for inordinate influence of the sociotechnical imaginaries [^13] Jasanoff and Kim [109] define sociotechnical imaginaries as “collectively held, institutionally stabilized, and publicly performed visions of desirable futures, animated by shared understandings of forms of social life and social order attainable through, and supportive of, advances in science and technology”. of profit-driven corporations over for example academic researchers [150]. This can thwart research that may not be immediately monetizable or challenges the status quo of power concentration, and complements the “regulatory capture” by bigger tech companies [136, 185, 19]. As Mitra [150] asks: “Whose sociotechnical imaginaries are granted normative status and what myriad of radically alternative futures are we overlooking?” Narratives of the inevitability of these technologies that are hyped up to be both transformative forces for society and simultaneously posing existential risks for humanity (often purported by the same actors) only bolster their imagined importance to accumulate increasing global investments, including from governments. Researchers who care about sociotechnical impact and ecological sustainability are busy with enumerating the harms of rapidly emerging new AI technologies and chasing potential mitigations instead of having the full means to imagine and develop systems for social good. While industry practitioners can contribute to both identifying new research challenges grounded in real-world systems and practical methods to mitigate some of the risks of emerging technologies, it is imperative that we create avenues for increasing independent research, while preserving the benefits of various modes of industry-academia collaborations.
Even as the grounded risks from these technologies (such as those discussed here) gather consensus from academic communities and civil society, it can be difficult to create space for alternative ways of development that are perceived as “slowing down”. Critical research on sociotechnical harms of AI is also under risk when attempts are made to shift attention from concerns about real harms to marginalized people today to unsubstantiated imagined future concerns [78, 77]. Calls for regulations to address these imagined future harms [79] further detract from real progress and contribute to reinforcement of monopolistic powers of those who have already added these technologies to their arsenals. This has led some sociotechnical researchers in AI to explicitly draw attention to how these systems shift power (e.g., [113, 29, 147, 76]), and to prioritize research guided by alternative visions for sociotechnical futures grounded in universal emancipation and social justice [150]. It is thus important that access to investments to enable development is also available to those trying to not only mitigate existing systems’ harms, but also develop new avenues, including work on social good and new business models.
@@
As generative AI starts to accumulate the lion’s share of research investments, it may starve out other areas of information access research. Generative AI has had exciting but limited deployments in information access systems today. There are significant open challenges to making these models broadly useful, including but not limited to concerns of potential sociotechnical harms. There is a risk that if these challenges are not mitigated in spite of the extensive resources already invested on them at present, there may be calls for even larger investments in future prompted by the sunk cost fallacy. [^14] https://en.wikipedia.org/wiki/Sunk_cost#Fallacy_effect It would be astute for the IR community to consciously continue to invest in research on systems and applications that societies need beyond what existing AI technologies make plausible [187, 150].
@@
Mechanism: Pollution of research artefacts Risks to academic research from generative AI may also emerge through the applications of generative AI models in IR scholarship—e.g., for authoring scientific papers and peer reviewing. There is evidence that researchers in computational sciences are already leveraging these tools [133], sometimes with hilariously terrible outcomes [168]. While the use of language models for light editing may (eventually) fall within the norms of socially acceptable behavior in research, their application in scholarship does raise concerns of plagiarism and scientific inaccuracies. This is an area that currently has more questions than answers and the IR community would benefit from proactively considering potential implications of this trend on future IR research.
@@ 2.1.5 Consequence: Ecological impact
Another important consequence of generative AI is its impact on the environment. In this context it is important for us to consider the direct environmental cost of developing and deploying generative AI systems at scale as well as the potential impact of these technologies on the climate change discourse online.
@@
Mechanism: Persuasive advertising Generative AI may not only negatively impact the environment through increasing demand for natural resources and increasing generation of waste, but may also supercharge climate change disinformation [74, 7, 9, 60, 197, 48]. For example, the fossil-fuel industry may attempt to sway public opinion through advertising that leverages generative AI’s persuasion capabilities discussed in §2.1.2. Persuasive advertising may also be employed by other environmentally-unfriendly business models like fast-fashion [47]. While the direct ecological cost of generative AI justifiably garners lots of attention, its potential impact on related online discourse also deserves scrutiny.
@@ 2.2.1 Risks to society
Information access is a critical need of any democratic society and a necessary ingredient for social transformation [102, 173, 84, 49, 86]. It is also a social determinant of economic progress [245, 156] and health [152]. Disruptions to the information ecosystem bears potentially grave risks to most aspects of our social lives. A confluence of the pandemic [186, 211, 70], rising global conflicts [210, 214], and escalating climate catastrophes [165, 108, 174] are pushing the world towards precarious instability. Our information ecosystems are already struggling under the weight of misinformation and disinformation that in this critical moment is eroding public trust in online platforms, institutions, and each other. It is imperative that researchers and developers of information access systems prioritize safeguarding social interests and be vigilant in considering potential risks of disruption and ecosystem collapse when integrating generative AI technologies in the IR stack. This includes identifying the necessary conditions under which these technologies can be safely deployed and developing practical safeguards and alternatives.
The risks to society are not just from potential disruptions of the information ecosystem, but also from how these technologies simultaneously concentrate power away from those at the margins of society. As institutions that develop and operate these technologies are themselves beneficiaries of this concentration, we need democratic oversights. If technologies further exacerbate already worsening wealth and power inequities, this additionally may pose severe threats to democratic institutions and human rights. There is an opportunity cost of not re-imagining information access in light of sociotechnical ambitions of human emancipation, culture, and knowledge production, instead of being constrained solely by what these emerging technologies make plausible and the homogenized visions put forth by institutions who wield these technologies [150].
@@ 2.2.2 Risks to IR research
IR research can suffer from a confluence of different factors including the distancing of academic researchers from the data and compute they need to do their work and how narratives about the inevitability of AI technologies shapes what computational research gets funded. The concentration of access to the networks around these technologies in a subset of institutions shapes what is considered “foundational” or even “AI”. Research on generative AI should not be performed only in the context of corporate economic interests while academia is hollowed out and prevented from exploring radical new methods that challenge the status quo. This risk of homogenization of academic research agendas and the opportunity cost of not exploring more diverse approaches to online information access can have material consequences. Instead, the IR community must be empowered with both the space and the resources necessary to explore a diversity of these visions and critique dominant narratives. IR research should have a plurality of work, which includes work with access to industry to change current practices. However, we especially also need to ensure that not all IR research is simply an extension of industrial system development and risk the demise of fundamental research on alternative avenues.
@@ 2.2.3 Risks to environment
Information access provides one of the large scale application settings for generative AI. However, the impact of such wide-scale deployment of these technologies on the impending climate crisis should be a critical consideration. Climate costs pose substantial existential risks for ecosystems and people, in more direct ways than some other “existential risks” that lack adequate scientific basis but have nonetheless been popular discourse in some parts of the AI community. This means both choosing what to deploy, and investment in methods to mitigate negative impacts that build on existing environmental work. As we discussed in §2.1.5, these concerns include not just the ecological cost of developing and deploying generative AI technologies but also their impact on online discourse on societal priorities.
@@ 3.1 Evaluating the impact of generative IR applications
Evaluating the impact of generative IR applications requires methods, as do data-informed interventions to steer that impact. Creating an LLM-based demo has become exceedingly easy. Understanding the impact of a system when it gets used in real life contexts, and getting to a high quality experience for a wide variety of users, is much harder. Standards for impact assessment have not kept up a similar pace as tech developments. Klaaf points out the need to carefully consider the differences in value alignment of the goals of a system, and safety considations, harms and risks [119]. A wide range of online, offline, and human-assisted evaluations are possible -and necessary- to get a full sense of the impact of a system.
@@ 3.2 Threat identification, assessment, and modeling
When the emergence of a new technology or application becomes apparent, the assessment of whether this poses risks or opportunities within specific domains poses a challenge. Before development of a system, threats and opportunities can be identified. As Kapoor et al. [115] point out, it’s crucial not to evaluate the risks and impact of new systems in isolation, but rather in comparison with existing technologies. For example, the impact of usage of foundation models in search should be compared to existing web search. For this purpose, Kapoor et al., present an evaluation framework focus on marginal risks, applied to Open Foundation Models. Their framework is based on threat identification work from cybersecurity and consists of six steps necessary to demonstrate such marginal risk. These steps are: 1) threat identification, 2) evaluating existing risk absent open foundation models, 3) considering existing defenses absent open foundation models, 4) evidence of marginal risk of open foundation models, 5) ease of defending against new risks, and 6) outlining uncertainty and assumptions. Note that this framework does not set exact assessment criteria, but rather defines the steps to get to such evaluations.
In practical settings, this might mean having to select standards for the development process (e.g. emerging standards from organizations such as NIST [8] or ISO [1], company-specific standards such as Microsoft’s Responsible AI Standard v2 General Requirements [2], or following new (local) legal requirements). However, mapping out potential consequences and identifying mechanisms that introduce risks in the specific context of a system needs to go much further. How to disrupt potential negative mechanisms in order to mitigate those risks requires gauging a wide range of consumer-side impacts [66], but also wider societal impacts. That includes frameworks focused on worker consequences [76], or practical methods focused on reducing the (legal) risks of using certain types of copyrighted or restricted training data vs. expected performance gains [148].